# Better Estimation of the KL Divergence Between Language Models

### v2

@@ F.1 Variance Reduction Proof
$\operatorname*{\mathbb{E}}\left[\left\lVert\nabla\delta_{\textsc{rb}}^{N}-G_{N}\right\rVert^{2}\right]$ (55a)
$=\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{\prime}}\left[\left\lVert\sum_{n=1}^{N}\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}}\left[\delta_{\textsc{mc}}^{n}(\overline{\boldsymbol{Y}}_{\leq n})\underbrace{\left(\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{Y}}_{<n})+\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}_{<n})\right)}_{\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{Y}}_{\leq n})}-G_{N}\,\middle|\,\overline{\boldsymbol{Y}}_{<n}=\overline{\boldsymbol{Y}}^{\prime}_{<n}\right]\right\rVert^{2}\right]$ (55b)
(definition of $\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(definition of $\nabla\delta_{\textsc{rb}}^{N}$)}}\end{subarray}$
$=\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{\prime}}\left[\left\lVert\sum_{n=1}^{N}\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{n}}\left[\delta_{\textsc{mc}}^{n}(\overline{\boldsymbol{Y}}^{n}_{\leq n})\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{Y}}^{n}_{\leq n})-G_{N}\,\middle|\,\overline{\boldsymbol{Y}}^{n}_{<n}=\overline{\boldsymbol{Y}}^{\prime}_{<n}\right]\right\rVert^{2}\right]$ (55c)
@@
$=\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{\prime}}\left[\left\lVert\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{1}}\left[\cdots\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{N}}\left[\sum_{n=1}^{N}\delta_{\textsc{mc}}^{n}(\overline{\boldsymbol{Y}}^{n}_{\leq n})\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{Y}}^{n}_{\leq n})-G_{N}\,\middle|\,\overline{\boldsymbol{Y}}^{N}_{<N}=\overline{\boldsymbol{Y}}^{\prime}_{<N}\right]\cdots\,\middle|\,\overline{\boldsymbol{Y}}^{1}_{<1}=\overline{\boldsymbol{Y}}^{\prime}_{<1}\right]\right\rVert^{2}\right]$ (55d)
(linearity of $\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(linearity of expectation)}}\end{subarray}$
$\leq\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{\prime}}\left[\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{1}}\left[\cdots\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{N}}\left[\left\lVert\sum_{n=1}^{N}\delta_{\textsc{mc}}^{n}(\overline{\boldsymbol{Y}}^{n}_{\leq n})\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{Y}}^{n}_{\leq n})-G_{N}\right\rVert^{2}\,\middle|\,\overline{\boldsymbol{Y}}^{N}_{<N}=\overline{\boldsymbol{Y}}^{\prime}_{<N}\right]\cdots\,\middle|\,\overline{\boldsymbol{Y}}^{1}_{<1}=\overline{\boldsymbol{Y}}^{\prime}_{<1}\right]\right]$ (55e)
(Jensen’s $\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(Jensen's inequality)}}\end{subarray}$
$=\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{1}}\left[\cdots\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{N}}\left[\left\lVert\sum_{n=1}^{N}\delta_{\textsc{mc}}^{n}(\overline{\boldsymbol{Y}}^{n}_{\leq n})\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{Y}}^{n}_{\leq n})-G_{N}\right\rVert^{2}\right]\right]$ (55f)
(law of total $\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(law of total expectation)}}\end{subarray}$
$=\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}}\left[\left\lVert\sum_{n=1}^{N}\delta_{\textsc{mc}}^{n}(\overline{\boldsymbol{Y}}_{\leq n})\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{Y}}_{\leq n})-G_{N}\right\rVert^{2}\right]$ (55g)
@@
$\leq\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}}\left[\left\lVert\sum_{n=1}^{N}\delta_{\textsc{mc}}^{n}(\overline{\boldsymbol{Y}}_{\leq n})\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{\prime}}\left[\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{Y}}_{\leq N}^{\prime})\mid\overline{\boldsymbol{Y}}^{\prime}_{\leq n}=\overline{\boldsymbol{Y}}_{\leq n}\right]-G_{N}\right\rVert^{2}\right]$ (55h)
[…]
@@
$=\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}}\left[\left\lVert\sum_{n=1}^{N}\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{\prime}}\left[\delta_{\textsc{mc}}^{n}(\overline{\boldsymbol{Y}}^{\prime}_{\leq n})\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{Y}}_{\leq N}^{\prime})-G_{N}\,\middle|\,\overline{\boldsymbol{Y}}^{\prime}_{\leq n}=\overline{\boldsymbol{Y}}_{\leq n}\right]\right\rVert^{2}\right]$ (55i)
@@
(linearity of $\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(linearity of expectation)}}\end{subarray}$
@@
$=\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}}\left[\left\lVert\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{\prime 1}}\left[\cdots\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{\prime N}}\left[\sum_{n=1}^{N}\delta_{\textsc{mc}}^{n}(\overline{\boldsymbol{Y}}^{\prime n}_{\leq n})\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{Y}}_{\leq N}^{\prime n})-G_{N}\,\middle|\,\boldsymbol{Y}^{\prime N}_{\leq N}=\boldsymbol{Y}_{\leq N}\right]\cdots\,\middle|\,\boldsymbol{Y}^{\prime 1}_{\leq 1}=\overline{\boldsymbol{Y}}_{\leq 1}\right]\right\rVert^{2}\right]$ (55j)
@@
(linearity of $\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(linearity of expectation)}}\end{subarray}$
@@
$\leq\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}}\left[\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{\prime 1}}\left[\cdots\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{\prime N}}\left[\left\lVert\sum_{n=1}^{N}\delta_{\textsc{mc}}^{n}(\overline{\boldsymbol{Y}}^{\prime n}_{\leq n})\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{Y}}_{\leq N}^{\prime n})-G_{N}\right\rVert^{2}\,\middle|\,\overline{\boldsymbol{Y}}^{\prime N}_{\leq N}=\overline{\boldsymbol{Y}}_{\leq N}\right]\cdots\,\middle|\,\overline{\boldsymbol{Y}}^{\prime 1}_{\leq 1}=\overline{\boldsymbol{Y}}_{\leq 1}\right]\right]$ (55k)
@@
(Jensen’s $\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(Jensen's inequality)}}\end{subarray}$
@@
$=\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{\prime 1}}\left[\cdots\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{\prime N}}\left[\left\lVert\sum_{n=1}^{N}\delta_{\textsc{mc}}^{n}(\overline{\boldsymbol{Y}}^{\prime n}_{\leq n})\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{Y}}_{\leq N}^{\prime n})-G_{N}\right\rVert^{2}\right]\right]$ (55l)
@@
(law of total $\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(law of total expectation)}}\end{subarray}$
@@
$=\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}^{\prime}}\left[\left\lVert\sum_{n=1}^{N}\delta_{\textsc{mc}}^{n}(\overline{\boldsymbol{Y}}^{\prime}_{\leq n})\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{Y}}_{\leq N}^{\prime})-G_{N}\right\rVert^{2}\right]$ (55m)
[…]
$=\operatorname*{\mathbb{E}}_{\overline{\boldsymbol{Y}}}\left[\left\lVert\nabla\delta_{\textsc{mc}}^{N}(\overline{\boldsymbol{Y}})-G_{N}\right\rVert^{2}\right].$ (55n)
@@
(definition of $\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(definition of $\nabla\delta_{\textsc{mc}}^{N}$)}}\end{subarray}$
See 8
@@
$\operatorname*{\mathbb{E}}\left[\left\lVert\nabla\delta_{\textsc{rb}}-G\right\rVert^{2}\right]$ $=\operatorname*{\mathbb{E}}\left[\left\lVert\lim_{N\rightarrow\infty}\nabla\delta_{\textsc{rb}}^{N}-G_{N}\right\rVert^{2}\right]$ (definition of $\nabla\delta_{\textsc{rb}}$) (56a)
$=\lim_{N\rightarrow\infty}\operatorname*{\mathbb{E}}\left[\left\lVert\nabla\delta_{\textsc{rb}}^{N}-G_{N}\right\rVert^{2}\right]$ (dominated convergence theorem, $\mathop{\mathrm{Var}}\left[{\normalcolor\nabla\delta_{\textsc{rb}}}\right]<\infty$) (56b)
$\leq\lim_{N\rightarrow\infty}\operatorname*{\mathbb{E}}\left[\left\lVert\nabla\delta_{\textsc{mc}}^{N}-G_{N}\right\rVert^{2}\right]$ (Lemma 13) (56c)
$=\operatorname*{\mathbb{E}}\left[\left\lVert\lim_{N\rightarrow\infty}\nabla\delta_{\textsc{mc}}^{N}-G_{N}\right\rVert^{2}\right]$ (dominated convergence theorem, $\mathop{\mathrm{Var}}\left[{\normalcolor\nabla\delta_{\textsc{mc}}}\right]<\infty$) (56d)
$=\operatorname*{\mathbb{E}}\left[\left\lVert\nabla\delta_{\textsc{mc}}-G\right\rVert^{2}\right].$ (Lemma 11) (56e)
F.2 A Note on Rao–Blackwellizing KL in Trust Region Algorithms
The conventional Monte Carlo estimator of $\textsc{kl}(p_{\boldsymbol{\theta}}\mid\mid q)$ used in the PPO algorithm in open-sourced RLHF libraries, e.g., (von Werra et al., 2020; Havrilla et al., 2023), is as follows:
$\delta_{\textsc{mc}}^{\textsc{ppo}}=\frac{1}{M}\sum_{m=1}^{M}\frac{p_{\boldsymbol{\theta}}(\boldsymbol{Y}^{(m)})}{p_{\boldsymbol{\theta}_{\text{old}}}(\boldsymbol{Y}^{(m)})}\log\frac{p_{\boldsymbol{\theta}_{\text{old}}}(\boldsymbol{Y}^{(m)})}{q(\boldsymbol{Y}^{(m)})},$ (57)

### v3

@@ D.1 Variance-Reduction Proof
$\operatorname*{\mathbb{E}}\left[\left\lVert\boldsymbol{\delta}_{\textsc{rb}}^{(N)}-\boldsymbol{G}^{N}\right\rVert^{2}\right]$ (53a)
$=\operatorname*{\mathbb{E}}_{{\overline{\boldsymbol{Y}}}^{\prime}}\left[\left\lVert\sum_{n=1}^{N}\operatorname*{\mathbb{E}}_{{\overline{\boldsymbol{Y}}}}\left[\mu_{\textsc{mc}}^{n}({\overline{\boldsymbol{Y}}}_{\leq n})\underbrace{\left(\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}({\overline{\boldsymbol{Y}}}_{<n})+\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}({\overline{Y}}_{n}\mid{\overline{\boldsymbol{Y}}}_{<n})\right)}_{\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}({\overline{\boldsymbol{Y}}}_{\leq n})}-\boldsymbol{G}^{N}\,\middle|\,{\overline{\boldsymbol{Y}}}_{<n}={\overline{\boldsymbol{Y}}}^{\prime}_{<n}\right]\right\rVert^{2}\right]$ (53b)
$\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(definition of $\boldsymbol{\delta}_{\textsc{rb}}^{(N)}$)}}\end{subarray}$
$=\operatorname*{\mathbb{E}}_{{\overline{\boldsymbol{Y}}}^{\prime}}\left[\left\lVert\sum_{n=1}^{N}\operatorname*{\mathbb{E}}_{{\overline{\boldsymbol{Y}}}^{n}}\left[\mu_{\textsc{mc}}^{n}({\overline{\boldsymbol{Y}}}^{n}_{\leq n})\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}({\overline{\boldsymbol{Y}}}^{n}_{\leq n})-\boldsymbol{G}^{N}\,\middle|\,{\overline{\boldsymbol{Y}}}^{n}_{<n}={\overline{\boldsymbol{Y}}}^{\prime}_{<n}\right]\right\rVert^{2}\right]$ (53c)
@@
$=\operatorname*{\mathbb{E}}_{{\overline{\boldsymbol{Y}}}^{\prime}}\left[\left\lVert\operatorname*{\mathbb{E}}_{{\overline{\boldsymbol{Y}}}^{1}}\left[\cdots\operatorname*{\mathbb{E}}_{{\overline{\boldsymbol{Y}}}^{N}}\left[\sum_{n=1}^{N}\mu_{\textsc{mc}}^{n}({\overline{\boldsymbol{Y}}}^{n}_{\leq n})\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}({\overline{\boldsymbol{Y}}}^{n}_{\leq n})-\boldsymbol{G}^{N}\,\middle|\,{\overline{\boldsymbol{Y}}}^{N}_{<N}={\overline{\boldsymbol{Y}}}^{\prime}_{<N}\right]\cdots\,\middle|\,{\overline{\boldsymbol{Y}}}^{1}_{<1}={\overline{\boldsymbol{Y}}}^{\prime}_{<1}\right]\right\rVert^{2}\right]$ (53d)
$\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(linearity of expectation)}}\end{subarray}$
$\leq\operatorname*{\mathbb{E}}_{{\overline{\boldsymbol{Y}}}^{\prime}}\left[\operatorname*{\mathbb{E}}_{{\overline{\boldsymbol{Y}}}^{1}}\left[\cdots\operatorname*{\mathbb{E}}_{{\overline{\boldsymbol{Y}}}^{N}}\left[\left\lVert\sum_{n=1}^{N}\mu_{\textsc{mc}}^{n}({\overline{\boldsymbol{Y}}}^{n}_{\leq n})\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}({\overline{\boldsymbol{Y}}}^{n}_{\leq n})-\boldsymbol{G}^{N}\right\rVert^{2}\,\middle|\,{\overline{\boldsymbol{Y}}}^{N}_{<N}={\overline{\boldsymbol{Y}}}^{\prime}_{<N}\right]\cdots\,\middle|\,{\overline{\boldsymbol{Y}}}^{1}_{<1}={\overline{\boldsymbol{Y}}}^{\prime}_{<1}\right]\right]$ (53e)
$\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(Jensen's inequality)}}\end{subarray}$
$=\operatorname*{\mathbb{E}}_{{\overline{\boldsymbol{Y}}}^{1}}\left[\cdots\operatorname*{\mathbb{E}}_{{\overline{\boldsymbol{Y}}}^{N}}\left[\left\lVert\sum_{n=1}^{N}\mu_{\textsc{mc}}^{n}({\overline{\boldsymbol{Y}}}^{n}_{\leq n})\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}({\overline{\boldsymbol{Y}}}^{n}_{\leq n})-\boldsymbol{G}^{N}\right\rVert^{2}\right]\right]$ (53f)
$\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(law of total expectation)}}\end{subarray}$
$=\operatorname*{\mathbb{E}}_{{\overline{\boldsymbol{Y}}}}\left[\left\lVert\sum_{n=1}^{N}\mu_{\textsc{mc}}^{n}({\overline{\boldsymbol{Y}}}_{\leq n})\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}({\overline{\boldsymbol{Y}}}_{\leq n})-\boldsymbol{G}^{N}\right\rVert^{2}\right]$ (53g)
$\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(${\overline{\boldsymbol{Y}}}^{1},\ldots,{\overline{\boldsymbol{Y}}}^{N}$ are i.i.d.)}}\end{subarray}$
$\leq\operatorname*{\mathbb{E}}_{{\overline{\boldsymbol{Y}}}}\left[\left\lVert\sum_{n=1}^{N}\mu_{\textsc{mc}}^{n}({\overline{\boldsymbol{Y}}}_{\leq n})\operatorname*{\mathbb{E}}_{{\overline{\boldsymbol{Y}}}^{\prime}}\left[\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}({\overline{\boldsymbol{Y}}}_{\leq N}^{\prime})\mid{\overline{\boldsymbol{Y}}}^{\prime}_{\leq n}={\overline{\boldsymbol{Y}}}_{\leq n}\right]-\boldsymbol{G}^{N}\right\rVert^{2}\right]$ (53h)
@@
$=\operatorname*{\mathbb{E}}_{{\overline{\boldsymbol{Y}}}}\left[\left\lVert\sum_{n=1}^{N}\operatorname*{\mathbb{E}}_{{\overline{\boldsymbol{Y}}}^{\prime}}\left[\mu_{\textsc{mc}}^{n}({\overline{\boldsymbol{Y}}}^{\prime}_{\leq n})\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}({\overline{\boldsymbol{Y}}}_{\leq N}^{\prime})-\boldsymbol{G}^{N}\,\middle|\,{\overline{\boldsymbol{Y}}}^{\prime}_{\leq n}={\overline{\boldsymbol{Y}}}_{\leq n}\right]\right\rVert^{2}\right]$ (53i)
@@
$\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(linearity of expectation)}}\end{subarray}$
@@
$=\operatorname*{\mathbb{E}}_{{\overline{\boldsymbol{Y}}}}\left[\left\lVert\operatorname*{\mathbb{E}}_{{\overline{\boldsymbol{Y}}}^{\prime 1}}\left[\cdots\operatorname*{\mathbb{E}}_{{\overline{\boldsymbol{Y}}}^{\prime N}}\left[\sum_{n=1}^{N}\mu_{\textsc{mc}}^{n}({\overline{\boldsymbol{Y}}}^{\prime n}_{\leq n})\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}({\overline{\boldsymbol{Y}}}_{\leq N}^{\prime n})-\boldsymbol{G}^{N}\,\middle|\,\boldsymbol{Y}^{\prime N}_{\leq N}=\boldsymbol{Y}_{\leq N}\right]\cdots\,\middle|\,\boldsymbol{Y}^{\prime 1}_{\leq 1}={\overline{\boldsymbol{Y}}}_{\leq 1}\right]\right\rVert^{2}\right]$ (53j)
@@
$\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(linearity of expectation)}}\end{subarray}$
@@
$\leq\operatorname*{\mathbb{E}}_{{\overline{\boldsymbol{Y}}}}\left[\operatorname*{\mathbb{E}}_{{\overline{\boldsymbol{Y}}}^{\prime 1}}\left[\cdots\operatorname*{\mathbb{E}}_{{\overline{\boldsymbol{Y}}}^{\prime N}}\left[\left\lVert\sum_{n=1}^{N}\mu_{\textsc{mc}}^{n}({\overline{\boldsymbol{Y}}}^{\prime n}_{\leq n})\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}({\overline{\boldsymbol{Y}}}_{\leq N}^{\prime n})-\boldsymbol{G}^{N}\right\rVert^{2}\,\middle|\,{\overline{\boldsymbol{Y}}}^{\prime N}_{\leq N}={\overline{\boldsymbol{Y}}}_{\leq N}\right]\cdots\,\middle|\,{\overline{\boldsymbol{Y}}}^{\prime 1}_{\leq 1}={\overline{\boldsymbol{Y}}}_{\leq 1}\right]\right]$ (53k)
@@
$\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(Jensen's inequality)}}\end{subarray}$
@@
$=\operatorname*{\mathbb{E}}_{{\overline{\boldsymbol{Y}}}^{\prime 1}}\left[\cdots\operatorname*{\mathbb{E}}_{{\overline{\boldsymbol{Y}}}^{\prime N}}\left[\left\lVert\sum_{n=1}^{N}\mu_{\textsc{mc}}^{n}({\overline{\boldsymbol{Y}}}^{\prime n}_{\leq n})\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}({\overline{\boldsymbol{Y}}}_{\leq N}^{\prime n})-\boldsymbol{G}^{N}\right\rVert^{2}\right]\right]$ (53l)
@@
$\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(law of total expectation)}}\end{subarray}$
@@
$=\operatorname*{\mathbb{E}}_{{\overline{\boldsymbol{Y}}}^{\prime}}\left[\left\lVert\sum_{n=1}^{N}\mu_{\textsc{mc}}^{n}({\overline{\boldsymbol{Y}}}^{\prime}_{\leq n})\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}({\overline{\boldsymbol{Y}}}_{\leq N}^{\prime})-\boldsymbol{G}^{N}\right\rVert^{2}\right]$ (53m)
@@
$\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(${\overline{\boldsymbol{Y}}}^{\prime 1},\cdots,{\overline{\boldsymbol{Y}}}^{\prime N}$ are i.i.d.)}}\end{subarray}$
$=\operatorname*{\mathbb{E}}_{{\overline{\boldsymbol{Y}}}}\left[\left\lVert\boldsymbol{\delta}_{\textsc{mc}}^{(N)}({\overline{\boldsymbol{Y}}})-\boldsymbol{G}^{N}\right\rVert^{2}\right].$ (53n)
$\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(definition of $\boldsymbol{\delta}_{\textsc{mc}}^{(N)}$)}}\end{subarray}$
See 5
@@
$\operatorname*{\mathbb{E}}\left[\left\lVert\boldsymbol{\delta}_{\textsc{rb}}-\boldsymbol{G}\right\rVert^{2}\right]$ $=\operatorname*{\mathbb{E}}\left[\left\lVert\lim_{N\rightarrow\infty}\boldsymbol{\delta}_{\textsc{rb}}^{(N)}-\boldsymbol{G}^{N}\right\rVert^{2}\right]$ (54a)
$\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(definition of $\boldsymbol{\delta}_{\textsc{rb}}$)}}\end{subarray}$
$=\lim_{N\rightarrow\infty}\operatorname*{\mathbb{E}}\left[\left\lVert\boldsymbol{\delta}_{\textsc{rb}}^{(N)}-\boldsymbol{G}^{N}\right\rVert^{2}\right]$ (54b)
$\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(dominated convergence theorem, $\mathop{\mathrm{Var}}\left[{\normalcolor\boldsymbol{\delta}_{\textsc{rb}}}\right]<\infty$)}}\end{subarray}$
$\leq\lim_{N\rightarrow\infty}\operatorname*{\mathbb{E}}\left[\left\lVert\boldsymbol{\delta}_{\textsc{mc}}^{(N)}-\boldsymbol{G}^{N}\right\rVert^{2}\right]$ (54c)
[…]
$=\operatorname*{\mathbb{E}}\left[\left\lVert\lim_{N\rightarrow\infty}\boldsymbol{\delta}_{\textsc{mc}}^{(N)}-\boldsymbol{G}^{N}\right\rVert^{2}\right]$ (54d)
$\begin{subarray}{c}\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(dominated convergence theorem, $\mathop{\mathrm{Var}}\left[{\normalcolor\boldsymbol{\delta}_{\textsc{mc}}}\right]<\infty$)}}\end{subarray}$
$=\operatorname*{\mathbb{E}}\left[\left\lVert\boldsymbol{\delta}_{\textsc{mc}}-\boldsymbol{G}\right\rVert^{2}\right].$ (54e)
[…]
D.2 A Note on Rao–Blackwellizing KL in Trust-Region Algorithms
The conventional Monte Carlo estimator of $\textsc{kl}(p_{\boldsymbol{\theta}}\mid\mid q)$ used in the PPO algorithm in open-sourced RLHF libraries, e.g., [40, 13], is as follows:
$\mu_{\textsc{mc}}^{\textsc{ppo}}=\frac{1}{M}\sum_{m=1}^{M}\frac{p_{\boldsymbol{\theta}}(\boldsymbol{Y}^{(m)})}{p_{\boldsymbol{\theta}_{\text{old}}}(\boldsymbol{Y}^{(m)})}\log\frac{p_{\boldsymbol{\theta}_{\text{old}}}(\boldsymbol{Y}^{(m)})}{q(\boldsymbol{Y}^{(m)})},$ (55)